OpenStudy (anonymous):

4x squared+ 16=-24x Solve by completing the square

OpenStudy (lgbasallote):

the first step in completing the square is to divide everything by the coefficient of x^2...so what is the coefficient of x^2?

OpenStudy (anonymous):

No idea.

OpenStudy (lgbasallote):

what is the number on the left side of x^2?

OpenStudy (anonymous):

4

OpenStudy (lgbasallote):

that's right! now divide every term by 4

OpenStudy (anonymous):

Okay, so x^2+4=-6x?

OpenStudy (lgbasallote):

yup :D now...transpose all the x on one side and all the constants on the other

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

all the x on one side...all the numbers on the other...like x^2 -x = 1 or something..can you do that?

OpenStudy (anonymous):

So like x^2=4=-4 so x^2=-8? -4 -4

OpenStudy (anonymous):

Oops, x^2=-10?

OpenStudy (lgbasallote):

no no...it's x^2 + 4 = -6x remember? you cannot combine 4 and 6x

OpenStudy (anonymous):

So it's x^2=6x-4?

OpenStudy (lgbasallote):

right..but bring -6x to the other side..put all the x on one side remember?

OpenStudy (anonymous):

x^2=-6x-4 wait, can you combine 6x and x^2? +6x +6x

OpenStudy (lgbasallote):

no. you cannot.. so your equation becomes..x^2 + 6x = -4 right? now here comes the good part...divide the coefficient of x (which is 6) by 2 then square it

OpenStudy (anonymous):

So x^2=6x=-4 6x/2=3x 3x^2=9x?

OpenStudy (lgbasallote):

no no...just the coefficient...just 6

OpenStudy (lgbasallote):

so it becomes 9

OpenStudy (anonymous):

Okay, but where does the coefficient come from?

OpenStudy (anonymous):

how do I get the 6?

OpenStudy (lgbasallote):

6x is the middle term right? the coefficient is the number on the left of the variable. the variable is x so the number on the left is 6. ok?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Aren't there 2 possible answers, 9 & something else?

OpenStudy (lgbasallote):

9 is not yet the answer... you just have x^2 + 6x +9 = -4 then via law of equality because you added 9 on the left side..you need to add 9 on the right side so... x^2 + 6x + 9 = -4 + 9 good?

OpenStudy (anonymous):

yeah, so x^2+6x+9=5?

OpenStudy (lgbasallote):

yes :DDD now factor out x^2 + 6x + 9

OpenStudy (anonymous):

Define factor out please, sorry I'm being so slow, I've been sick for a while, and as a result I have no idea what's going on, plus I usually get slightly ignorant when I'm sick, I assure you, I'm usually much better at Math.

OpenStudy (lgbasallote):

x^2 + 6x + 9 is a perfect square trinomial..by factor out i mean change it to square of a binomial

OpenStudy (anonymous):

I know what you're talking about, but my brain really doesn't want to work right now.

OpenStudy (lgbasallote):

Hint: what was the number you squared a while ago?

OpenStudy (anonymous):

6?

OpenStudy (lgbasallote):

no...what is the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

right! so the square of binomial is (x+3)^2 plug it in... (x_3)^2 = 5 what do you think comes next?

OpenStudy (anonymous):

4891

OpenStudy (anonymous):

Lol, sorry, just being dumb. :P

OpenStudy (lgbasallote):

4891? no...you take the square root of both sides. :)))

OpenStudy (anonymous):

So (x-3)^2=5 x-3= Isn't the square root of 5 something like 2.2?

OpenStudy (lgbasallote):

leave it as sqrt (5) but it's +/- remember now transpose -3 and you're all done

OpenStudy (anonymous):

How do I transpose -3?

OpenStudy (anonymous):

x-3=sqrt(5) x=sqrt(5)+3 +3 +3